Question 1 (1 point)

Saved

Ultrasonic sound that travels at 1530 m/s in ocean water is bounced off the ocean floor and returns to the ship above. If the round trip takes 3 s, the depth of the water is

Question 1 options:

2295 m.

1530 m.

3060 m.

4590 m.

Question 2 (1 point)

Saved

A bass fiddle is louder than a harp because of its

Question 2 options:

sounding board.

thicker strings.

lower pitch.

all of the above

none of the above

Question 3 (1 point)

Saved

A 60 vibration per second wave travels 30 m in 1 s. Its frequency is

Question 3 options:

30 Hz and it travels at 60 m/s.

60 Hz and it travels at 30 m/s.

neither

Question 4 (1 point)

Saved

Two tuning forks produce sounds of wavelengths 3.4 m and 3.3 m. What beat frequency is produced?

Question 4 options:

2.0 Hz

1.0 Hz

0.1 Hz

4.0 Hz

3.0 Hz

Question 5 (1 point)

Saved

Resonance can be viewed as forced vibration with the

Question 5 options:

maximum amount of energy input.

minimum beat frequency.

matching of wave amplitudes.

least amount of energy input.

matching of constructive and destructive interference.

Question 6 (1 point)

Saved

The explanation for all types of refraction involves a change in

Question 6 options:

speed.

period.

frequency.

all of the above

none of the above

Question 7 (1 point)

Saved

A skipper on a boat notices wave crests passing the anchor chain every 5 s. The skipper estimates the distance between crests is 15 m. What is the speed of the water waves?

Question 7 options:

15 m/s

3 m/s

5 m/s

not enough information

Question 8 (1 point)

Saved

Sound travels fastest in

Question 8 options:

ice

water vapor

water

Question 9 (1 point)

Saved

A compression is a squash and a rarefaction is a

Question 9 options:

reverse tension.

crunch.

stretch.

Question 10 (1 point)

Saved

For light, a red shift indicates the light source moves

Question 10 options:

toward you.

at right angles to you.

away from you.

all of the above

none of the above

Question 11 (1 point)

Saved

On days when the air nearest the ground is colder than air above, sound waves

Question 11 options:

tend to be refracted upward.

travel without refraction.

tend to be refracted downward.

Question 12 (1 point)

Saved

Sound waves can interfere with one another so that no sound is produced.

Question 12 options:

True

False

Question 13 (1 point)

Saved

How many vibrations per second are associated with a 101-MHz radio wave?

Question 13 options:

more than 101,000,00

101,000,000

less than 101,000,000

Question 14 (1 point)

Saved

The pitch of a musical sound depends on the sound wave's

Question 14 options:

amplitude.

frequency.

wavelength.

speed.

all of the above

Question 15 (1 point)

Saved

A standing wave occurs when

Question 15 options:

a wave reflects upon itself.

the speed of the wave is zero or near zero.

two waves overlap.

the amplitude of a wave exceeds its wavelength.

Question 16 (1 point)

Saved

A wiggle in time is a

Question 16 options:

wave.

vibration.

both

neither

Question 17 (1 point)

Saved

The loudness of a musical sound is a measure of the sound wave's

Question 17 options:

wavelength.

amplitude.

frequency.

speed.

all of the above

Question 18 (1 point)

Saved

The vibrations of a transverse wave move in a direction

Question 18 options:

at right angles to the direction of wave travel.

opposite to wave travel.

along the direction of wave travel.

Question 19 (1 point)

Saved

The Doppler effect is characteristic of

Question 19 options:

sound waves.

light waves.

water waves.

all of the above

none of the above

Question 20 (1 point)

Saved

A wave travels an average distance of 1 m in 1 s with a frequency of 1 Hz. Its amplitude is

Question 20 options:

not enough information

1 m.

more than 1 m.

less than
Answer #1

Q1)-

If we assume the distance between ship an d ocean floor = x than-

2x = 1530 * t where t =3 sec

2x = 1530 * 3

x = 2295 m

Q2)-

Sounding Board

Q3)- Frequency means vibration per second i.e 60 Hz

And Velocity of the wave will be 30m/s

Q4)- Beat Frq = |f2 - f1|

f2 = c / = 3*108 / 3.3

f2 = 0.9090*108 hz

f1 = c / = 3*108 / 3.4

f1 = 0.88235 *108 Hz

Beat Freq = | f2 -f1|

= 0.02673 *108 Hz

Q5)-

matching of wave amplitudes.

Q6)-

The change of direction of light as it passes from one medium to another is associated with a change in velocity and wavelength.

Q7)-

s = vt

15 = v*5

V = 3 m/s

Now rest of the questions u have to ask in the another uploaded question set.

Thankyou !!! Have Fun

Please Rate High.